OpenStudy (anonymous):

Solve the equation using the quadratic formula. 2x2 + 12x + 3 = 0

OpenStudy (anonymous):

Can you identify a, b, and c?

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (nory):

a=the coefficient of 2x2 b=the coefficient of 12x c=the coefficient of 3 Do you know what a coefficient is?

OpenStudy (anonymous):

coefficient is basically the number seen before its term

OpenStudy (nory):

Yes. So write out all of the coefficients as a, b, and c as I described.

OpenStudy (anonymous):

i am not the one who asked the question :P

OpenStudy (anonymous):

I know how to the problem until the end. here \[-12\pm \sqrt{120}\div4\]